OpenStudy (anonymous):

Determine the mean,median,mode, and midrange of the set of data. Round your answer to the nearest tenth: 4,6,10,12,10,9,365,40,37,8.

OpenStudy (anonymous):

For median, place them in order 4,6,8,9, 10, 10,12,37,40,365 ^ ^ | | there are two numbers in the middle. Add them up and divide by 2 10 + 10 = 20 and 20/2 = 10 10 is the median

OpenStudy (anonymous):

Mode is the number that occurs the most. There are more 10's therefore the mode is 10

OpenStudy (anonymous):

For midrange, take the biggest and smallest number, add them up and divide by 2 365+4 = 369 369/2 = 164.5

OpenStudy (anonymous):

For mean, add them all up and divide by 10

OpenStudy (anonymous):

Thank you so much for your help:)